# Eyes Closed, Safety On: Protecting Multimodal LLMs via Image-to-Text Transformation

### v1

@@ 0.D.0.1 Training with model-generated data
has become an essential research problem in both computer vision (e.g., He et al. [25] for image classification, GeoDiffusion [8, 19, 38, 31] for object detection [24, 29] and also StableRep [54] for contrastive learning [5, 39] and masked image modeling [6, 66]) and natural language processing (e.g., SELF [40] for instruction tuning and mistake analysis [7] for safety alignment) and vision-language modeling [28, 21], thanks to the remarkable progress of AIGC. ECSO also belongs to this direction, but different from previous works, we focus on inheriting the intrinsic safety mechanism of pre-aligned LLMs to safeguard MLLMs to the greatest extend, as in Sec. 4.

### v3

@@ 0.D.0.1 Training with model-generated data
has become an essential research problem in both computer vision (e.g., He et al. [26] for image classification, GeoDiffusion [8, 20, 41, 32, 62, 19] for object detection [25, 30] and StableRep [57] for contrastive learning [5, 43] and masked image modeling [6, 71]) and natural language processing (e.g., SELF [44] for instruction tuning and mistake analysis [7] for safety alignment) and vision-language modeling [29, 22], thanks to the remarkable progress of AIGC. ECSO also belongs to this direction. However, different from previous works, we focus on inheriting the intrinsic safety mechanism of pre-aligned LLMs to safeguard MLLMs to the greatest extent, as in Sec. 4.